calculate number of moles of O-atom in126 amu ofHNO3/

Grade:12th pass

1 Answers

Arun
5 years ago

1 amu = 1.66 x 10-24 grams

126 amu = 126 x 1.66 x 10-24

              = 209.16 x 10-24 grams

1 mole HNO3 = 63.01 g

Hence 209.16 x 10-24 grams = 3.32 x 10-24 moles

1 mole = 6.023 x 1023 molecules

hence 3.32 x 10-24 moles = 19.99 x 10-23  molecules

1 molecule of HNOcontains 5 atoms.

Hence, 19.99 x 10-23  molecules = 99.96 x 10-23 atoms

Therefore, 126 amu of HNO3 contains 99.96 x 10-23 atoms.